The recording of airtraffic came back in operation at about: 11:50 A.M.
There has been no recording of air trafic since 7.December.
The recording of airtraffic went down today at about: 7:04 P.M.
Stream Cam3 and Black and white cam has been repaired and is in now operation.
Updated the software in the Teradel streaming devices (cam2 and cam3). Hopefully this will make the clock info in the screen more reliable.
Finished cleaning up and packing. Go home.
Students and teachers go home from Hessdalen. Science Camp is finished.
Students and teachers from Greaker vgs attend Hessdalen. They will run Science Camp.
Roger Eide from Inspiria Science Center, Bjørn Gitle Hauge and Anna-Lena Kjøniksen from Østfold College
attend Hessdalen. They will run The Science Camp.
Erling Strand go to Hessdalen, preparing Science Camp.
The recording of airtraffic came back in operation at about: 2:10 P.M.
There has been no recording of air trafic since 21.June.
Erling Strand travel back home from Hessdalen. It was a good tour: Great weather and some observations. The northern measurements station
for the french scientists was down. Problems with the batteries. The southern station was ok. Did restart the PC's in the blue box.
There are some trouble with the alarm PC. It goes very often down.
The windows on cam2 and cam3 has been cleaned today.
The "big black dirt" on cam3 is gone.
Erling Strand (me) travel to Hessdalen for doing a guide tour, having observation nights with PH friends.
The measurement stations will also be checked.
The recording of airtraffic went down today at about: 2:10 P.M.
The recording of airtraffic came back in operation at about: 10:50 A.M.
There has been no recording of air trafic since 2.June.
The recording of airtraffic went down today at about: 4:00 P.M.
There is a new page about interesting pictures from 2013.
There is also one report.
There is a new page about interesting pictures from 2014.
These pictures are take automatically. These (preliminary) two pictures are from
Gerhard Gröschel automatics station at Peder farm.
The alarm-PC in the station (Blue Box) has been repaired, and it will now start sending alarms to
soso-alarmpage again. The cam3 and geiger counter has been repaired.
The cam3 had a bad connector, so it has sometimes been working. Thanks to Jan Moen for fixing this.
Today, a new page decribing an observation from 1976